### v1

@@ VII-G Tranformers
The integration of transformer architectures, such as bidirectional encoder representations from Transformer (BERT) and connectionist temporal classification (CTC) among others [219], into the realm of KC diagnosis and segmentation represents a cutting-edge approach that holds significant promise for advancing medical imaging analysis. Drawing insights from several studies, such as those referenced, the potential applications of transformers in this domain are evident. One of the primary advantages of transformers lies in their ability to capture long-range dependencies within the data, making them particularly suitable for tasks involving complex relationships and patterns, as often encountered in medical imaging. The referenced studies showcase various transformer-based models tailored specifically for KC diagnosis and segmentation.
@@
Furthermore, the study by Wang et al. [221] showcases the synergy between transformers and convolutional neural networks (CNNs) for renal parenchymal tumors diagnosis. By stacking transformers and CNNs, the model efficiently captures both long-range dependencies and low-level spatial details, enhancing its performance in identifying and diagnosing kidney lesions. Moreover, transformer-based methods, such as TA-UNet3+ proposed by Hu [222], introduce innovative attention mechanisms and encoder-decoder architectures to improve kidney tumor segmentation accuracy. Similarly, Qian et al. [223] present a hybrid network combining nnU-Net and Swin Transformer, emphasizing the critical role of precise delineation and localization in KC diagnosis and treatment planning. Additionally, transformers exhibit promise in histopathological image analysis for subtype classification of renal cell carcinoma (RCC). Gao et al. [224] demonstrate the effectiveness of instance-based vision transformers in subtyping papillary RCC, providing valuable insights for histopathological diagnosis and personalized treatment strategies.
VII-H Quantum Computing
Quantum computing represents a nascent and burgeoning domain poised to transform our approach to computational challenges, including those within ML, with the potential to revolutionize various aspects of knowledge discovery [225]. Despite being at an early developmental stage, quantum computing harbors the promise of significantly enhancing the efficacy of TL models in the diagnosis of diseases like KC [226]. Employing quantum computing in TL for KC diagnosis entails harnessing quantum algorithms to tackle tasks that would otherwise be impractical with classical computers, notably optimization and linear algebra computations. These algorithms offer avenues to bolster TL model performance by curtailing computational overheads and enhancing model accuracy [227]. Additionally, quantum computing holds promise in mitigating classical computing constraints encountered in medical imaging, such as the demand for substantial computational resources and memory for processing voluminous datasets [228, 229]. By leveraging quantum computing, the duration required for training and inference can be slashed, fostering swifter diagnoses and improved patient outcomes [230, 231].
While quantum computing holds immense promise for revolutionizing cancer diagnosis using transfer and DL techniques, significant advancements in quantum algorithms, hardware capabilities, and data availability are necessary to realize its full potential in improving the diagnosis and treatment of KC and other types of cancer [232]. Collaborative efforts between quantum physicists, computer scientists, and medical researchers are essential to overcome these challenges and pave the way for the future of quantum-enhanced healthcare technologies [233].
VII-I Large Language Models and Generative Chatbots
@@
Large language models (LLMs) and generative chatbots hold significant potential for transforming KC diagnosis in several ways [234, 235]. By integrating vast medical libraries and research data, LLMs can provide healthcare professionals with up-to-date information, diagnostic suggestions, and treatment options based on the latest studies and clinical guidelines [236, 237]. Moving on, generative chatbots can interpret patients’ descriptions of their symptoms in natural language, aiding in the early detection of KC by flagging potential warning signs and suggesting further diagnostic tests [238, 239, 240]. Typically, these models can deliver personalized information to patients about KC, including explanations of their condition, possible treatment pathways, and what to expect during each treatment option, improving patient understanding and engagement in their care plan [241, 242, 243].
@@
Besides, LLMs can analyze vast amounts of data from EHRs to identify patterns or risk factors associated with KC, potentially uncovering new insights into its diagnosis and progression [244, 245]. With training, LLMs can assist radiologists by providing a second opinion on imaging studies, such as CT scans and MRIs, identifying areas that may warrant a closer look for signs of KC [246, 247, 248]. Additionally, chatbots can handle routine inquiries and patient triaging, allowing medical staff to focus on more complex tasks and patient care, thereby increasing the efficiency of clinical workflows [249, 250]. LLMs can integrate and analyze information from various sources, including genomic data, radiographic images, and clinical notes, offering a holistic view of a patient’s health status and aiding in the accurate diagnosis of KC [251, 252]. Moreover, generative chatbots can facilitate remote monitoring of patient’s symptoms and treatment responses, providing timely advice and enabling telehealth consultations with specialists [253, 254].
VIII Conclusion
The application of TL and adaptive domain techniques presents a promising avenue for enhancing the precision of KC diagnoses. TL enables the adaptation of models trained on extensive datasets to new, smaller datasets where labeled data may be scarce. Meanwhile, adaptive domain techniques allow models to adjust to shifts in data distribution over time, a critical factor for maintaining accuracy in the dynamic landscape of medical diagnostics. By employing TL, models pre-trained on comprehensive medical image datasets can undergo fine-tuning with smaller sets of kidney images for more precise KC diagnosis. This approach leverages the pre-training phase’s acquired relevant features, proving advantageous when faced with limited labeled data availability. The adaptive domain further refines model accuracy as data distribution evolves, a key aspect in KC diagnostics where new data continuously emerges. Ensuring the model’s adaptability to these changes is essential for sustaining diagnostic precision.
Our research delves into the utilization of TL in identifying KC, highlighting its potential to mitigate computational demands compared to conventional DL models, circumvent annotated data shortages, and handle dataset variability across training and testing phases. The paper’s organizational section lays out a detailed discourse on various concepts and classifications. It deep-dives into the literature surrounding TL-driven KC detection, offering an extensive classification of contributions to the field. Subsequently, it proposes several general solutions to the challenges encountered in KC detection, applicable to both traditional and TL-based models. Drawing from the reviewed literature, it advocates for future methodological advancements aimed at precise and accurate KC detection, underscoring the importance of harnessing TL’s full spectrum of capabilities.
Despite the successes of TL and DA in KC detection, numerous challenges warrant further exploration. One such challenge is the homogeneous nature of many DA models, which assume consistent feature distances between source and target domains—an assumption not always valid. Some studies have ventured into knowledge transfer without this prerequisite, utilizing datasets to enhance task support. Moreover, as highlighted, the future is likely to see heightened focus on heterogeneous adaptive domain approaches until their optimization is achieved, thereby advancing KC detection systems.
Data availability
Data will be made available on request.
Conflict of Interest

### v2

@@ 7.9 Tranformers and Large language models for KC
The integration of transformer architectures, such as bidirectional encoder representations from Transformer (BERT) and connectionist temporal classification (CTC) among others [194, 195], into the realm of KC diagnosis and segmentation represents a cutting-edge approach that holds significant promise for advancing medical imaging analysis. Drawing insights from several studies, such as those referenced, the potential applications of transformers in this domain are evident. One of the primary advantages of transformers lies in their ability to capture long-range dependencies within the data, making them particularly suitable for tasks involving complex relationships and patterns, as often encountered in medical imaging. The referenced studies showcase various transformer-based models tailored specifically for KC diagnosis and segmentation.
@@
Furthermore, the study by Wang et al. [197] showcases the synergy between transformers and CNNs for renal parenchymal tumors diagnosis. By stacking transformers and CNNs, the model efficiently captures both long-range dependencies and low-level spatial details, enhancing its performance in identifying and diagnosing kidney lesions. Moreover, transformer-based methods, such as TA-UNet3+ proposed by Hu [198], introduce innovative attention mechanisms and encoder-decoder architectures to improve kidney tumor segmentation accuracy. Similarly, Qian et al. [199] present a hybrid network combining nnU-Net and Swin Transformer, emphasizing the critical role of precise delineation and localization in KC diagnosis and treatment planning. Additionally, transformers exhibit promise in histopathological image analysis for subtype classification of RCC. Gao et al. [200] demonstrate the effectiveness of instance-based vision transformers in subtyping papillary RCC, providing valuable insights for histopathological diagnosis and personalized treatment strategies.
[…]
@@
LLM and generative chatbots hold significant potential for transforming KC diagnosis in several ways [201, 202]. By integrating vast medical libraries and research data, large language model (LLM) can provide healthcare professionals with up-to-date information, diagnostic suggestions, and treatment options based on the latest studies and clinical guidelines [203, 204]. Moving on, generative chatbots can interpret patients’ descriptions of their symptoms in natural language, aiding in the early detection of KC by flagging potential warning signs and suggesting further diagnostic tests [205, 206, 207]. Typically, these models can deliver personalized information to patients about KC, including explanations of their condition, possible treatment pathways, and what to expect during each treatment option, improving patient understanding and engagement in their care plan [208, 209, 210].
@@
Besides, LLM can analyze vast amounts of data from EHRs to identify patterns or risk factors associated with KC, potentially uncovering new insights into its diagnosis and progression [211, 212]. With training, LLM can assist radiologists by providing a second opinion on imaging studies, such as CT scans and MRIs, identifying areas that may warrant a closer look for signs of KC [213, 214, 215]. Additionally, chatbots can handle routine inquiries and patient triaging, allowing medical staff to focus on more complex tasks and patient care, thereby increasing the efficiency of clinical workflows [216, 217]. LLM can integrate and analyze information from various sources, including genomic data, radiographic images, and clinical notes, offering a holistic view of a patient’s health status and aiding in the accurate diagnosis of KC [218, 219]. Moreover, generative chatbots can facilitate remote monitoring of patient’s symptoms and treatment responses, providing timely advice and enabling telehealth consultations with specialists [220, 221].
8 Conclusion
The integration of TL and DA techniques marks a pivotal advancement in enhancing the accuracy, efficiency, and robustness of KC diagnosis. TL plays a critical role in overcoming the limitations of traditional DL approaches by enabling models pre-trained on large, diverse datasets to adapt effectively to smaller, specialized datasets, where annotated data is often scarce—a persistent challenge in medical diagnostics. This knowledge transfer not only reduces the computational burden of training complex models from scratch but also significantly improves diagnostic precision, even when fine-tuning with limited kidney-specific data. Complementing this, DA techniques address the critical issue of data distribution shifts, ensuring that diagnostic models remain reliable and accurate amidst the dynamic and evolving nature of medical imaging data. Through a systematic and comprehensive analysis of TL-based frameworks for KC detection, this study provides a rigorous evaluation of key contributions, identifies pressing challenges, and proposes actionable solutions to push the boundaries of the field. It underscores the transformative potential of TL in tackling major hurdles, including the computational complexity of DL, the scarcity of annotated medical data, and the variability between training and testing datasets, thereby enhancing diagnostic reliability and precision. Despite these advancements, significant challenges persist. Many DA models rely on the assumption of homogeneity in feature distributions between source and target domains—an assumption that often fails to hold in real-world medical scenarios, where data heterogeneity is the norm. Recent research has begun to explore heterogeneous DA approaches, which relax these assumptions and open new avenues for innovation. Looking ahead, future research should prioritize the optimization of these heterogeneous DA methods to improve their applicability in KC detection, fostering the development of more robust, generalizable, and scalable diagnostic systems. Furthermore, efforts should focus on integrating these techniques into clinical workflows, ensuring seamless translation from research to practice. By addressing these challenges, the synergy of TL and DA has the potential to revolutionize precision medicine, driving transformative changes in clinical practice, improving patient outcomes, and advancing the frontiers of oncology research. However, some limitations of this review must be acknowledged. The studies reviewed primarily focus on a narrow subset of methods, often omitting emerging techniques in the broader TL and DA fields that may offer significant improvements. Additionally, there is a heavy reliance on studies with limited clinical validation, which could impact the generalizability of the findings to real-world medical environments. Some studies also show methodological inconsistencies in how TL and DA techniques are applied, making direct comparisons challenging. Furthermore, the review has been constrained by the availability of high-quality, peer-reviewed papers, limiting the inclusion of non-published, yet potentially valuable, research. Future reviews should strive to include a broader range of studies, especially those that explore new, diverse, and interdisciplinary approaches to TL and DA in medical imaging.
Declarations
@@ 8 Conclusion
  • Funding: The authors declare that there is no fund for this project
  • Data availability: There is no data available for this project
  • Ethics approval and consent to participate: Not Applicable.